Solve the equations
2x-5=y
2y=6x-9
Not really sure how to set it up to even start to solve it.
Please help and Thak you in advance.

Found 2 solutions by mananth, rfer:
Answer by mananth(16949)

2x-5=y
2y=6x-9
..
rearrange the terms such that the variable are on one side and the numbers on the other side.
2x-5=y
add +5 to both sides
2x-5+5=y+5
2x=y+5
add -y to both sides
2x-y=y+5+y
2x-y=5.......1
..
the other equation will be
2y-6x =-9............2
..
multiply equation 1 by 3 and to equation 2
..
6x-3y+(2y-6x)=15-9
6x-3y+2y-6x=6
-y=6
y=-6
..
plug the value of y in 2x-y=5
2x-(-6)=5
2x+6=5
2x=-1
x=-1/2

Answer by rfer(16322)

2x-5=y
2y=6x-9
2(2x-5)=6x-9
4x-10=6x-9
-2x=1
x=-.5
------------
2*-.5-5=y
-1-5=y
y=-6
-------------
2*-.5-5=-6
-1=-1